CVP Analysis Test Acctg 48 V2 Mona Company incurs $80,000 of annual fixed costs in manufacturing and selling a product that it sells for $16 per unit The variable costs of manufacturing and selling the product are $9 per unit a) The contribution margin per unit for the product is b) The breakeven point in units is c) The breakeven point in dollars is If Mona Company has a 30% income tax rate, and its management wants to earn an annual after-tax income of $35,000. Determine the dollar sales volume company must achieve to reach this target income level d) e) Determine the amount of after-tax income that Mona Company would eam from $420,000 sales volume
Answer #1
Req a:
Selling price 16
Less: variable cost 9
CM per unit 7
Req b:
Break even units: fixed cost / CMm per unit
80000/7 = 11429 units
Req c:
CM ratio = CM per unit/ Selling price *100
7/16 *100 = 43.75 per unit
Break even in $ = Fixed cost/ CM ratio
80000 / 43.75% = $ 182857
Req d:
Target after tax income = 35000
Pre tax income = 35000 /70% = 50000
Desired contribution = 80000+50000 = 130000
target sales in $ = Desired contribution / CM ratio
130000 /43.75% = $ 297143
Req e:
Salles 420000
Cm ratio 43.75%
Contribution 183750
Less: Fixed cosot 80000
Pre tax income 103750
Less: Tax @ 30% 31125
After tax income 72625